OpenStudy (anonymous):

if a root of the equation x^3-4x^2+2x+4=0 is equal to the sum of the other two roots, find the roots

OpenStudy (anonymous):

to t the sum of the other two roots?

OpenStudy (anonymous):

opps sorry, that's a typo

OpenStudy (anonymous):

uh i have the three roots

OpenStudy (anonymous):

how did you find them?

OpenStudy (anonymous):

-.7320508076 2 2.7320508076

OpenStudy (anonymous):

hm.. how did you do that?

OpenStudy (anonymous):

well i saw that (x-2) could be taken out of the first three terms

OpenStudy (anonymous):

sorry, I'm not sure how you took it out, did you factorise the x out?

OpenStudy (anonymous):

https://www.dpmms.cam.ac.uk/~wtg10/cubic.html

OpenStudy (anonymous):

All of the knowledge on solving cubic equations.

OpenStudy (anonymous):

aw thanks! But we haven't been taught this yet. Is there another way of doing this question?

OpenStudy (anonymous):

Well it gets confusing after cubic equations.. i normally try to find the quadratic equation in there and then take it out. and the rest will factor itself out... but not really 100% sure

OpenStudy (anonymous):

I don't know either. I tried making the roots a, b, (a+b), found the product and sum of roots and tried to solve it from there but it wouldn't work

OpenStudy (anonymous):

hi duplicity do you know synthetic division? use it here

OpenStudy (anonymous):

or just divide the eq by 2 to get x^3-4x^2+2x+4=0 (x-2)(x^2-2x-2)=0 now we know the root x1=2 you can find the other two roots by using the quadratic formula

OpenStudy (anonymous):

hi mark, how did you divide the eq by 2 to get (x-2)(x^2-2x-2)=0?

OpenStudy (anonymous):

im sorry divide by x-2

OpenStudy (anonymous):

ah I see! How did you know to divide by that?

OpenStudy (anonymous):

i use synthetic division, are you familiar with that?

OpenStudy (anonymous):

sorry, nope. but what is it?

OpenStudy (anonymous):

it is one of good tool for finding roots of polynomials by first guessing

OpenStudy (anonymous):

oh! is it where you sub in factors of the remainder to check if it makes a zero?

OpenStudy (anonymous):

yes it is try this one http://en.wikipedia.org/wiki/Synthetic_division

OpenStudy (anonymous):

or try this http://mathworld.wolfram.com/SyntheticDivision.html

OpenStudy (anonymous):

Ok, I got it! It like another version of polynomial division. Thank you!!

OpenStudy (anonymous):

yes it is ..lol ok yw good luck now